OpenStudy (anonymous):

P(A)=0.4 P(B)=0.3 and P(A and B)=0.6 find P(A or B)

OpenStudy (raden):

use the formula : P(A or B) = P(A) + P(B) - P(A and B)

OpenStudy (anonymous):

So 0.4+0.3-0.6 equals P(A or B)?

OpenStudy (cwrw238):

P(A ∪ B) = P(A) + P(B) - P(A ∩ B))

OpenStudy (cwrw238):

right

OpenStudy (zarkon):

"P(A)=0.4 P(B)=0.3 and P(A and B)=0.6 " this is not possible

OpenStudy (zarkon):

the event (A and B) is a subset of A and a subset of B. thus \[P(A\text{ and }B)\le P(A)=.4\] \[P(A\text{ and }B)\le P(B)=.3\] thus \[P(A\text{ and }B)\le.3\] and so it can't equal .6

OpenStudy (anonymous):

ty:)))